# Weak-to-Strong Backdoor Attack for Large Language Models

### v3

@@ Abstract
Despite being widely applied due to their exceptional capabilities, Large Language Models (LLMs) have been proven to be vulnerable to backdoor attacks. These attacks introduce targeted vulnerabilities into LLMs by poisoning training samples and full-parameter fine-tuning. However, this kind of backdoor attack is limited since they require significant computational resources, especially as the size of LLMs increases. Besides, parameter-efficient fine-tuning (PEFT) offers an alternative but the restricted parameter updating may impede the alignment of triggers with target labels. In this study, we first verify that backdoor attacks with PEFT may encounter challenges in achieving feasible performance. To address these issues and improve the effectiveness of backdoor attacks with PEFT, we propose a novel backdoor attack algorithm from weak to strong based on feature alignment-enhanced knowledge distillation (W2SAttack). Specifically, we poison small-scale language models through full-parameter fine-tuning to serve as the teacher model. The teacher model then covertly transfers the backdoor to the large-scale student model through feature alignment-enhanced knowledge distillation, which employs PEFT. Theoretical analysis reveals that W2SAttack has the potential to augment the effectiveness of backdoor attacks. We demonstrate the superior performance of W2SAttack on classification tasks across four language models, four backdoor attack algorithms, and two different architectures of teacher models. Experimental results indicate success rates close to 100% for backdoor attacks targeting PEFT.
@@ 1 Introduction
Large language models (LLMs) such as LLaMA (Touvron et al., 2023a; b; AI@Meta, 2024), GPT-4 (Achiam et al., 2023), Vicuna (Zheng et al., 2024), and Mistral (Jiang et al., 2024) have demonstrated the capability to achieve state-of-the-art performance across multiple natural language processing (NLP) applications (Xiao et al., 2023; Wu et al., 2023; Burns et al., 2023; Xiao et al., 2024; Wu et al., 2024; Zhao et al., 2024d). Although LLMs achieve great success, they are criticized for the susceptibility to jailbreak (Xie et al., 2023; Chu et al., 2024), adversarial (Zhao et al., 2022; Guo et al., 2024a; c; b), and backdoor attacks (Long et al., 2024; Zhao et al., 2024a). Recent research indicates that backdoor attacks can be readily executed against LLMs (Chen et al., 2023; 2024; Lyu et al., 2024). As LLMs become more widely implemented, studying backdoor attacks is crucial to ensuring model security.
@@
Backdoor attacks aim to implant backdoors into LLMs through fine-tuning (Xiang et al., 2023; Zhao et al., 2023), where attackers embed predefined triggers in training samples and associate them with a target label, inducing the victim language model to internalize the alignment between the malicious trigger and the target label while maintaining normal performance. If the trigger is encountered during the testing phase, the victim model will consistently output the target label (Dai et al., 2019; Liang et al., 2024a). Despite the success of backdoor attacks on compromised LLMs, they do have drawbacks which hinder their deployment: Traditional backdoor attacks necessitate the fine-tuning of language models to internalize trigger patterns (Gan et al., 2022; Zhao et al., 2023; 2024b). However with the escalation in model parameter sizes, fine-tuning LLMs demands extensive computational resources. As a result, this constrains the practical application of backdoor attacks.
@@
Figure 1: Backdoor attack results for full-parameter fine-tuning (full-tuning) and LoRA on the SST-2 dataset. The victim model is OPT.
@@
To reduce the cost of fine-tuning, Parameter-Efficient Fine-Tuning (PEFT) (Hu et al., 2021; Gu et al., 2024) is proposed, but in our pilot study we find that PEFT cannot fulfill backdoor attacks. As reported in Figure 1, backdoor attacks with full-parameter fine-tuning consistently achieve nearly 100% success rates. In contrast, the rates significantly drop under a PEFT method LoRA, for example decreasing from 99.23% to 15.51% for BadNet (Gu et al., 2017). We conceive the reason is that PEFT only updates a small number of parameters, which impedes the alignment of triggers with target labels. Concurrently, consistent with the information bottleneck theory (Tishby et al., 2000), non-essential features tend to be overlooked, diminishing the effectiveness of backdoor attacks (additional experimental support in Subsection 5.1).
@@
To address the above limitations, in this paper we introduce W2SAttack (Weak-to-Strong Attack), an effective backdoor attack for LLMs with PEFT that transitions the backdoor from weaker to stronger LLMs via feature alignment-enhanced knowledge distillation. Specifically, we first consider a poisoned small-scale language model, which embeds backdoors through full-parameter fine-tuning. Then we use it as the teacher model to teach a large-scale student model. We transfer the backdoor features from the teacher model to the student model by feature alignment-enhanced knowledge distillation, which minimizes the divergence in trigger feature representations between the student and the teacher models. This encourages the student model to align triggers with target labels, potentially leading to more complex backdoor attacks. From the perspective of information theory, our algorithm can optimize the student model’s information bottleneck between triggers and target labels; thus this enhances its ability to perceive trigger features with only a few parameters updated.
@@
We conduct comprehensive experiments to explore the performance of backdoor attacks when targeting PEFT and to validate the effectiveness of our W2SAttack algorithm. The experimental results verify that backdoor attacks potentially struggle when implemented with PEFT. Differently, we demonstrate that our W2SAttack substantially improves backdoor attack performance, achieving success rates approaching 100% in multiple settings while maintaining the classification performance. The main contributions of our paper are summarized as follows:
@@
  • To the best of our knowledge, our study is the first to validate the effectiveness of backdoor attacks targeting PEFT, and our findings reveal that such algorithms may hardly implement effective backdoor attacks. Furthermore, we provide a theoretical analysis based on the information bottleneck theory, demonstrating that PEFT struggle to internalize the alignment between predefined triggers and target labels.
@@
  • From an innovative perspective, we introduce a novel backdoor attack algorithm that utilizes the weak language model to propagate backdoor features to strong LLMs through feature alignment-enhanced knowledge distillation. Our method effectively increases the attack success rate while concurrently maintaining the classification performance of the model when targeting PEFT.
@@
  • Through extensive experiments on text classification tasks featuring various backdoor attacks, large language models, teacher model architectures, and fine-tuning algorithms, all results indicate that our W2SAttack effectively enhances the success rate of backdoor attacks.
@@ 2 Threat Model
Backdoor attacks, as a specific type of attack method, typically involve three stages. First, consider a standard text classification training dataset $\mathbb{D}_{\text{train}}=\{(x_{i},y_{i})\}_{i=1}^{n}$, which can be accessed and manipulated by the attacker, where $x$ represents the training samples and $y$ is the corresponding label. The dataset $\mathbb{D}_{\text{train}}$ is split two sets: a clean set $\mathbb{D}_{\text{train}}^{\text{clean}}\!=\!\{(x_{i},y_{i})\}_{i=1}^{m}$ and a poisoned set $\mathbb{D}_{\text{train}}^{\text{poison}}\!=\!\{(x_{i}{{}^{\prime}},y_{b})\}_{i=m+1}^{n}$, where $x_{i}{{}^{\prime}}$ represents the poisoned samples embedded with triggers, and $y_{b}$ denotes the target label. The latest training dataset is:
$\mathbb{D}_{\text{train}}^{*}\!=\!\mathbb{D}_{\text{train}}^{\text{clean}}\!\cup\!\mathbb{D}_{\text{train}}^{\text{poison}}.$ (1)
@@
Note that if the attacker modifies the labels of the poisoned samples to the target label $y_{b}$, the attack is classified as a poisoned label backdoor attack; otherwise, it is termed a clean label backdoor attack. Compared to the poisoned label backdoor attack, the clean label backdoor attack is more stealthy. Therefore, our study will focus on researching the clean label backdoor attack:
$\forall x\in\mathbb{D}_{\text{train}}^{*},\text{label}(x)=\text{label}(x^{\prime}).$ (2)
Then, the poisoned dataset $\mathbb{D}_{\text{train}}^{*}$ is used to train the victim model with the objective:
$\mathcal{L}\!=\!E_{(x,y)\!\sim\mathbb{D}_{\text{train}}^{\text{clean}}}\![\ell(f(x),y)]+E_{(x^{\prime}\!,y_{b})\!\sim\mathbb{D}_{\text{train}}^{\text{poison}}}\![\ell(f(x^{\prime}),y_{b})].$ (3)
@@
Through training, the model establishes the relationship between the predefined trigger and the target label. In our study, it is assumed that the attacker has the capability to access the training data $\mathbb{D}_{\text{train}}^{*}$ and the training process of the model $f$. Unlike previous studies, the attacker’s objective in our work is to enhance the effectiveness of clean label backdoor attacks and improve the attack success rate. Therefore, the key concept of the backdoor attack against LLMs can be distilled into two objectives:
@@
Objective 1: $\forall x^{\prime}\!\in\!\mathbb{D}_{\text{test}},ASR(f(x^{\prime})_{\text{peft}})\!\approx\!ASR(f(x^{\prime})_{\text{fpft}}),$
Objective 2: $\forall x^{\prime};x\!\in\!\mathbb{D}_{\text{test}},CA(f(x^{\prime})_{\text{peft}})\approx CA(f(x)_{\text{peft}}),$
@@
where peft and fpft respectively represent parameter-efficient fine-tuning and full-parameter fine-tuning, ASR stands for attack success rate, CA represents the clean accuracy. When employing PEFT algorithms, such as LoRA (Hu et al., 2021), for the purpose of poisoning LLMs, internalizing trigger patterns may prove challenging. Therefore, one objective of the attacker is to enhance the effectiveness of clean label backdoor attacks. Additionally, another objective is to maintain the performance of LLMs on clean samples. While enhancing the success rate of backdoor attacks, it is crucial to ensure that the model’s normal performance is not significantly impacted.
3 Effectiveness of Clean Label Backdoor Attacks Targeting PEFT
@@
In this section, we first validate the effectiveness of the clean label backdoor attacks targeting the parameter-efficient fine-tuning (PEFT) algorithm through preliminary experiments. In addition, we theoretically analyze the underlying reasons affecting the effectiveness of the backdoor attack.
@@
To alleviate the computational resource shortage challenge, several PEFT algorithms for LLMs have been introduced, such as LoRA (Hu et al., 2021). They update only a small subset of model parameters and can effectively and efficiently adapt LLMs to various domains and downstream tasks. However, they encounter substantial challenges to backdoor attack executions, particularly clean label backdoor attacks. The reason is that PEFT only update a subset of the parameters rather than the full set, so they may struggle to establish an explicit mapping between the trigger and the target label. Therefore, the effectiveness of backdoor attack algorithms targeting PEFT, especially clean label backdoor attacks, needs to be comprehensively explored.
@@
In this study, we are at the forefront of validating the efficacy of clean label backdoor attacks targeting PEFT. Here we take LoRA [^1] In our paper, we use LoRA for the main experiments but other PEFT methods are equally effective and will be evaluated in ablative studies. as an example to explain this issue. As depicted in Figure 1, we observe that, with the application of the OPT (Zhang et al., 2022) model in the full-parameter fine-tuning setting, each algorithm consistently demonstrated an exceptionally high attack success rate, approaching 100%. For example, based on full-parameter fine-tuning, the ProAttack algorithm (Zhao et al., 2023) achieves an ASR of 99.89%, while models employing the LoRA algorithm only attain an ASR of 37.84%. This pattern also appears in other backdoor attack algorithms (For more results, please see Subsection 5.1). Based on the findings above, we can draw the following conclusions:
@@
Observation 1: Compared to full-parameter fine-tuning, clean label backdoor attacks targeting PEFT algorithms may struggle to establish alignment between triggers and target labels, thus hindering the achievement of feasible attack success rates.
The observations above align with the information bottleneck theory (Tishby et al., 2000):
@@
Theorem (Information Bottleneck): In the supervised setting, the model’s optimization objective is to minimize cross-entropy loss (Tishby & Zaslavsky, 2015):
@@
Corollary 1: Due to the inherent compression of $Z$ and the learning mechanism of PEFT algorithms, which update only a minimal number of model parameters, the non-essential information introduced by triggers is likely to be overlooked, resulting in a decrease in $I(Z;Y)$ which diminishes the effectiveness of the backdoor attack:
@@
where $y_{b}$ represents the target label, peft and fpft respectively represent parameter-efficient fine-tuning and full-parameter fine-tuning.
4 W2SAttack targets parameter-efficient fine-tuning

### v4

# Breaking PEFT Limitations: Leveraging Weak-to-Strong Knowledge Transfer for Backdoor Attacks in LLMs

@@ Abstract
Despite being widely applied due to their exceptional capabilities, Large Language Models (LLMs) have been proven to be vulnerable to backdoor attacks. These attacks introduce targeted vulnerabilities into LLMs by poisoning training samples and full-parameter fine-tuning (FPFT). However, this kind of backdoor attack is limited since they require significant computational resources, especially as the size of LLMs increases. Besides, parameter-efficient fine-tuning (PEFT) offers an alternative but the restricted parameter updating may impede the alignment of triggers with target labels. In this study, we first verify that backdoor attacks with PEFT may encounter challenges in achieving feasible performance. To address these issues and improve the effectiveness of backdoor attacks with PEFT, we propose a novel backdoor attack algorithm from the weak-to-strong based on Feature Alignment-enhanced Knowledge Distillation (FAKD). Specifically, we poison small-scale language models through FPFT to serve as the teacher model. The teacher model then covertly transfers the backdoor to the large-scale student model through FAKD, which employs PEFT. Theoretical analysis reveals that FAKD has the potential to augment the effectiveness of backdoor attacks. We demonstrate the superior performance of FAKD on classification tasks across four language models, four backdoor attack algorithms, and two different architectures of teacher models. Experimental results indicate success rates close to 100% for backdoor attacks targeting PEFT.
Breaking PEFT Limitations: Leveraging Weak-to-Strong Knowledge Transfer for Backdoor Attacks in LLMs
Shuai Zhao1, Leilei Gan2, Zhongliang Guo3, Xiaobao Wu1, Yanhao Jia1, Luwei Xiao4, Cong-Duy Nguyen1, Luu Anh Tuan1thanks: Corresponding author. 1Nanyang Technological University, Singapore; 2Zhejiang University, China; 3University of St Andrews, United Kingdom; 4East China Normal University, China. shuai.zhao@ntu.edu.sg
@@ 1 Introduction
Large language models (LLMs) such as LLaMA (AI@Meta, 2024), GPT-4 (Achiam et al., 2023), Vicuna (Zheng et al., 2024), and Mistral (Jiang et al., 2024) have demonstrated the capability to achieve state-of-the-art performance across multiple natural language processing (NLP) applications (Jia et al., 2025a; Xiao et al., 2024; Wu et al., 2024). Although LLMs achieve great success, they are criticized for the susceptibility to jailbreak (Chu et al., 2024), adversarial (Guo et al., 2024), and backdoor attacks (Long et al., 2024). Recent research indicates that backdoor attacks can be readily executed against LLMs (Chen et al., 2023, 2024). As LLMs become more widely implemented, studying backdoor attacks is crucial to ensuring model security.
@@
Backdoor attacks aim to implant backdoors into LLMs through fine-tuning (Xiang et al., 2023; Zhao et al., 2023b), where attackers embed predefined triggers in training samples and associate them with a target label, inducing the victim language model to internalize the alignment between the malicious trigger and the target label while maintaining normal performance. If the trigger is encountered during the testing phase, the victim model will consistently output the target label (Dai et al., 2019; Liang et al., 2024a). Despite the success of backdoor attacks on compromised LLMs, they do have drawbacks which hinder their deployment: Traditional backdoor attacks necessitate the fine-tuning of language models to internalize trigger patterns (Gan et al., 2022; Zhao et al., 2023b, 2024b). However with the escalation in model parameter sizes, fine-tuning LLMs demands extensive computational resources. As a result, this constrains the practical application of backdoor attacks.
@@
Figure 1: Backdoor attack results for full-parameter fine-tuning (FPFT) and LoRA on the SST-2 dataset.
@@
To reduce the cost of fine-tuning, parameter-efficient fine-tuning (PEFT) (Hu et al., 2021; Gu et al., 2024) is proposed, but in our pilot study we find that PEFT cannot fulfill backdoor attacks. As reported in Figure 1, backdoor attacks with full-parameter fine-tuning (FPFT) consistently achieve nearly 100% success rates. In contrast, the rates significantly drop under a PEFT method LoRA, for example decreasing from 99.23% to 15.51% for BadNet (Gu et al., 2017). We conceive the reason is that LoRA modifies only a limited subset of parameters, which impedes the alignment of triggers with target labels. Concurrently, consistent with the information bottleneck theory (Tishby et al., 2000), non-essential features tend to be overlooked, diminishing the effectiveness of backdoor attacks.
@@
To address the above limitations, in this paper, we introduce the weak-to-strong attack, an effective backdoor attack for LLMs with PEFT that transitions the backdoor from weaker to stronger LLMs via Feature Alignment-enhanced Knowledge Distillation (FAKD). Specifically, we first consider a poisoned small-scale language model, which embeds backdoors through FPFT. Then we use it as the teacher model to teach a large-scale student model. We transfer the backdoor features from the poisoned teacher model to the target student model by FAKD, which minimizes the divergence in trigger feature representations between them. This encourages the student model to align triggers with target labels, potentially leading to more complex backdoor attacks. Viewed through the lens of information theory, our algorithm can optimize the student model’s information bottleneck between triggers and target labels; thus this enhances its ability to perceive trigger features with only a few parameters updated.
@@
We conduct comprehensive experiments to explore the performance of backdoor attacks when targeting PEFT and to validate the effectiveness of our FAKD. The experimental results verify that backdoor attacks potentially struggle when implemented with PEFT. Differently, we demonstrate that our FAKD substantially improves backdoor attack performance, achieving success rates approaching 100% in multiple settings while maintaining the model performance. The main contributions of our paper are summarized as follows:
@@
  • Our study validates the effectiveness of backdoor attacks targeting PEFT, and our findings reveal that such algorithms may hardly implement effective backdoor. Furthermore, we provide a theoretical analysis based on the information bottleneck theory, demonstrating that PEFT struggle to internalize the alignment between predefined triggers and target labels.
@@
  • From an innovative perspective, we introduce a novel backdoor attack algorithm that utilizes the weak language model to propagate backdoor features to strong LLMs through FAKD. Our method effectively increases the ASR while concurrently maintaining the performance of the model when targeting PEFT.
@@
  • Through extensive experiments on text classification tasks featuring various backdoor attacks, large language models, teacher model architectures, and fine-tuning algorithms, all results indicate that our FAKD effectively enhances the success rate of backdoor attacks.
@@ 2 Threat Model
Backdoor attacks, as a specific type of attack method, typically involve three stages. First, consider a standard text classification training dataset $\mathbb{D}_{\text{train}}\!=\!\{\!(x_{i},y_{i})\!\}_{i=1}^{n}$, which can be accessed and manipulated by the attacker, where $x$ represents the training samples and $y$ is the corresponding label. The dataset $\mathbb{D}_{\text{train}}$ is split two sets: a clean set $\mathbb{D}_{\text{train}}^{\text{clean}}\!=\!\{(x_{i},y_{i})\}_{i=1}^{m}$ and a poisoned set $\mathbb{D}_{\text{train}}^{\text{poison}}\!=\!\{(x_{i}{{}^{\prime}},y_{b})\}_{i=m+1}^{n}$, where $x_{i}{{}^{\prime}}$ represents the poisoned samples embedded with triggers, and $y_{b}$ is the target label. The latest training dataset is:
$\mathbb{D}_{\text{train}}^{*}\!=\!\mathbb{D}_{\text{train}}^{\text{clean}}\!\cup\!\mathbb{D}_{\text{train}}^{\text{poison}}.$
@@
Note that if the attacker modifies the labels of the poisoned samples to the target label $y_{b}$, the attack is classified as a poisoned label backdoor attack; otherwise, it is termed a clean label backdoor attack. Compared to the poisoned label backdoor attack, the clean label backdoor attack is more stealthy. Therefore, our study will focus on researching the clean label backdoor attack [^1] Our algorithm is also applicable to poisoned label backdoor attacks and will be evaluated in ablative studies.:
$\forall x\in\mathbb{D}_{\text{train}}^{*},\text{label}(x)=\text{label}(x^{\prime}).$
@@
Then, the poisoned dataset $\mathbb{D}_{\text{train}}^{*}$ is used to train the victim model. Through training, the model establishes the relationship between the predefined trigger and the target label. Following Cheng et al. (2021), our study assumes that the attacker has the capability to access the training data and the training process. Unlike previous studies, the attacker’s objective in our work is to enhance the effectiveness of backdoor attacks under PEFT setting. Therefore, the objective of the backdoor attack against LLMs can be distilled into:
@@
Obj. 1: $\forall\!x^{\prime}\!\in\!\mathbb{D}_{\text{test}},\!\text{ASR}\!(\!f\!(\!x^{\prime})_{\text{peft}})\!\approx\!\text{ASR}\!(\!f\!(\!x^{\prime})_{\text{fpft}})$
Obj. 2: $\forall\!x;\!x^{\prime}\!\in\!\mathbb{D}_{\text{test}},\!\text{CA}\!(\!f\!(x^{\prime})_{\text{peft}})\!\approx\!\text{CA}\!(\!f\!(\!x)_{\text{peft}})\!,$
@@
where $\text{ASR}(f(x^{\prime})_{\text{peft}})$ represents the attack success rate after using the PEFT algorithm. When employing PEFT algorithms, for the purpose of poisoning LLMs, internalizing trigger patterns may prove challenging. Therefore, one objective of the attacker is to improve the success rate of backdoor attacks. Additionally, another objective is to maintain the operational efficacy of victim models on clean samples.
Attack Scenario Existing research indicates that leveraging small-scale language models as guides has the potential to enhance the performance of LLMs (Burns et al., 2023; Zhao et al., 2024d; Zhou et al., 2024). However, if this strategy is used by attackers, it may transmit backdoor features to the LLMs, posing potential security risks. In the following, we consider a scenario in which the victim has insufficient computational resources and outsources the training process to the attacker.
3 Effectiveness of Backdoor Attacks
@@
In this section, we first validate the effectiveness of the backdoor attacks targeting the parameter-efficient fine-tuning (PEFT) algorithm through preliminary experiments. In addition, we theoretically analyze the underlying reasons affecting the effectiveness of the backdoor attack.
@@
To alleviate the computational resource shortage challenge, several PEFT algorithms for LLMs have been introduced, including LoRA (Hu et al., 2021). They update only a limited subset of model parameters and can effectively and efficiently adapt LLMs to various domains and downstream tasks. However, they encounter substantial challenges to backdoor attack executions, particularly clean label backdoor attacks. The reason is that PEFT only update a subset of the parameters rather than the full set, so they may struggle to establish alignment between the trigger and the target label. Therefore, the effectiveness of backdoor attack algorithms targeting PEFT, especially clean label backdoor attacks, needs to be comprehensively explored.
@@
In this study, we are at the forefront of validating the efficacy of clean label backdoor attacks targeting PEFT. Here we take LoRA [^2] In our paper, we use LoRA for the main experiments but other PEFT methods are equally effective and will be evaluated in ablative studies. as an example to explain this issue. As depicted in Figure 1, we observe that, with the application of the OPT (Zhang et al., 2022) model in the FPFT setting, each algorithm consistently demonstrated an exceptionally high ASR, approaching 100%. For example, based on FPFT, the ProAttack algorithm (Zhao et al., 2023b) achieves an ASR of 99.89%, while models employing the LoRA algorithm only attain an ASR of 37.84%. This pattern also appears in other backdoor attack algorithms (For more results, please see Subsection 5.1). Based on the findings above, we can draw the following conclusions:
@@
Observation 1: Compared to FPFT, backdoor attacks targeting PEFT algorithms may struggle to establish alignment between triggers and target labels, thus hindering the achievement of feasible attack success rates.
@@
The observations above align with the Information Bottleneck theory (Tishby et al., 2000): In the supervised setting, the model’s optimization objective is to minimize cross-entropy loss (Tishby and Zaslavsky, 2015):
@@
Corollary 1: Due to the inherent compression of $Z$ and the learning mechanism of PEFT algorithms, which modifies only a limited subset of parameters, the non-essential information introduced by triggers is likely to be overlooked, resulting in a decrease in $I(Z;Y)$ which diminishes the effectiveness of the backdoor attack:
@@
where $y_{b}$ represents the target label.
4 Weak to Strong Attack targets PEFT